# Automating High Quality RT Planning at Scale

### v1

@@ Abstract
Radiotherapy (RT) planning is complex, subjective, and time-intensive. Advances in artificial intelligence (AI) promise to improve its precision, efficiency, and consistency, but progress is often limited by the scarcity of large, standardized datasets. To address this, we introduce the Automated Iterative RT Planning (AIRTP) system, a scalable solution for generating high-quality treatment plans. This scalable solution is designed to generate substantial volumes of consistently high-quality treatment plans, overcoming a key obstacle in the advancement of AI-driven RT planning. Our AIRTP pipeline adheres to clinical guidelines and automates essential steps, including organ-at-risk (OAR) contouring, helper structure creation, beam setup, optimization, and plan quality improvement, using AI integrated with RT planning software like Varian’s Eclipse. Furthermore, a novel approach for determining optimization parameters to reproduce 3D dose distributions, i.e. a method to convert dose predictions to deliverable treatment plans constrained by machine limitations. A comparative analysis of plan quality reveals that our automated pipeline produces treatment plans of quality comparable to those generated manually, which traditionally require several hours of labor per plan. Committed to public research, the first data release of our AIRTP pipeline includes nine cohorts covering head-and-neck and lung cancer sites to support an AAPM 2025 challenge. This data set features more than 10 times the number of plans compared to the largest existing well-curated public data set to our best knowledge.
Repo: github.com/RiqiangGao/GDP-HMM_AAPMChallenge.
@@ 1 Introduction
Radiation therapy (RT) is a critical treatment modality used in approximately 50% of all cancer cases. Deep learning approaches have been incorporated into various stages of the RT planning process, including dose prediction [17, 6, 12], fluence prediction [22], leaf sequencing [11], and dose calculation [23]. Among recent advancements in RT, 3D dose prediction emerges as a particularly promising innovation. This approach directly estimates 3D dose distributions from inputs such as CT scans, PTV/OAR masks, and planning configurations, eliminating the need for labor-intensive optimization processes. Dose prediction serves multiple purposes, including supporting optimization objectives [6, 5], enabling quality assurance [13, 14], and functioning as a key component in fully AI-driven planning pipelines [11].
RT planning is a complex and subjective process involving multiple domain experts, making data curation from existing clinical plans highly labor-intensive due to human variability in contouring PTVs/OARs and planning practices. For instance, the number of OARs contoured and the quality of plans can vary significantly (see more in Sec. 2.1). Variations in radiotherapy planning introduce biases into training data, which are further exacerbated when data is collected from multiple cohorts. Additionally, most dose prediction studies face challenges such as a limited number of training patients and difficulties in quantitatively assessing quality metrics. Within the same institution, plan quality can also fluctuate over time due to the adoption of upgraded optimization technologies or updates to treatment protocols. These changes lead to inconsistencies in plan quality, further complicating the data used for AI training and evaluation.
AI-based organ segmentation has achieved significant advancements in both research and commercial applications. Varian Medical Systems’ Eclipse, a widely used platform for optimizing radiotherapy treatment plans for cancer patients, supports automation through its Eclipse Scripting API (ESAPI). ESAPI enables developers to access and manipulate patient data for planning purposes, streamlining traditionally manual tasks performed via the Eclipse Graphics User Interface (GUI). This functionality enhances workflow efficiency and ensures greater consistency in radiotherapy planning.
@@
  • We develop a method to replicate 3D dose distributions in both complex IMRT and VMAT scenarios using AIRTP, facilitating the evaluation of dose prediction impacts on deliverable dose planning.
  • We have released the largest cleaned planning dataset to date for an AAPM challenge [^2] The GDP-HMM is one of the selected challenges by the AAPM Working Group on Grand Challenges in 2025 (www.aapm.org/GrandChallenge). and beyond. This dataset: 1) contains over 10 times more plans than the existing curated public dataset [6], and 2) offers more practical settings to support diverse research explorations.
@@ 2.1 Challenges of Clinical Plans for AI Training
Additionally, some clinical plans were created more than five or ten years ago, when radiotherapy technology was less advanced. As a result, these plans may not meet the higher quality standards enabled by current technology. This highlights the need for re-planning RT treatments using the latest advancements while preserving the original clinical intent.
@@ 2.2 Automatic High Quality Planning at Scale
Step 2: Planning Script. There are three major modules in the planning script. Beam configuration includes setting up the fractions, beam geometries, and machine parameters. For each patient, we have both IMRT and VMAT setups following the same clinical goals. For the optimizer and dose calculation, we use TrueBeam with Photon Optimization version PO.18.0.1 and Dose Calculation AXB 18.0.1, respectively. To balance the time cost, the optimization convergence is set to ‘On‘ for VMAT.
@@
More detailed steps are outlined in Algorithm 1, along with some module implementations shown in Figures 3 and 4.
Input: Patient data with CT, PTV/OAR contours, helper structures, prescribed doses for PTVs; ScoreCard for plan evaluation; beam configurations; RapidPlan for a head-and-neck scenario. Output: A series of plans until the ScoreCard convergence. 1:Initialize plan based on RapidPlan or ScorePlan items 2:for $n$ = 1 to MAX iteration do 3:   Initialize objective buffer $D$ 4:   for $s$ in PTV and its Ring structures do 5:      Set the upper / lower point objectives based prescribed dose 6:      Use the same priorities as in the initial plan 7:   end for 8:   Calculate the ScoreCard for current plan 9:   for $s$ in ScoreCard OARs and Helpers do 10:      Calculate the DVH Cumulative Data $f_{s}(\cdot)$ 11:      for $v$ in volume points {1, 10, 20, 30, 40, 60, 80do 12:         Calculate the dose $d=f_{s}(v)$ 13:         Calculate the margin and objective dose and priority based on 3. 14:         Add the objective to objective buffer. 15:      end for 16:      for items in ScoreCard do 17:         Calculate the margin and objective dose and priority based on 4. 18:         Add the objective to objective buffer. 19:      end for 20:   end for 21:   optimize with the current objective buffer. 22:   update current plan. 23:end for
@@
(a) iterative high-quality planning
(b) translate 3D dose prediction to deliverable plan
3 Planning and Results
3.1 Planning Environment
3.2 Plan Quality Analysis
Fig. 6 demonstrates plan examples for lung cancer and head-and-neck cancer patients. Our AIRTP can generate arbitrarily large number of plans with varying parameters, given sufficient time and computing resources. As the first public release of our data, we include two plans per patient for the lung site and three plans per patient for the head-and-neck site. Additional examples can be found in Appendices B and C.
(a) Score improvement of IMRT plans.
(b) Score improvement of VMAT plans.
Figure 7: Quality improvement achieved through iterative refinement across all head-and-neck cohorts.
3.3 Comparison with Clinical Plans
3.4 Plan Reproducibility Given 3D Dose
(a) HaN: reference (solid) vs. reproduce (dash).
(b) Lung: reference (solid) vs. reproduce (dash).
3.5 Beyond the Topics of GDP-HMM Challenge
4 Discussion
As an initial step toward automating high-quality planning at scale, our work has several limitations compared to expert planning in clinical practice. First, clinical planning involves detailed patient-specific considerations, such as variations in the contouring of PTVs/OARs, auxiliary structures, and various dose constraints. In our pipeline, we focus primarily on clinical PTVs and prescribed doses as the key patient-specific factors, simplifying the process due to the lack of access to more detailed patient data. Second, different institutions often follow different protocols, and even within the same institution, planners may have individual preferences. Our pipeline does not currently account for such subjective variations. While these preferences are partially reflected in the scorecards, we use only one standardized scorecard per treatment site. A future direction is to collaborate with dosimetrists to develop multiple scorecards that better capture diverse clinical preferences and practices. Third, unlike more well-established research fields, our comparative analyses of iterative refinement and re-planning strategies are limited by the lack of directly relevant literature. Although we made efforts to compare the DVHs of AIRTP plans with clinically approved plans, this comparison may not be entirely fair, as the AIRTP and clinical plans follow different guidelines. Fourth, the summary and clinical impact analysis of the GDP-HMM challenge are beyond the scope of this paper and will be provided following the conclusion of the challenge.
Appendix
[…]
Appendix A Detailed Scorecard Examples
@@
Table 7 and 8 shows the details scores and final score from the scorecard [20] to evaluate plan qualities for IMRT and VMAT, respectively. The structures (from the automatic pipeline) too small or with too much overlap with PTVs have been removed. For each table, we have included the RT structure names (i.e., StructId), Metric Type, the input for the Metric Type, output value, and the score of the output. We compare the output and its value of the plans from the initial (without the iterative refining) and full AIRTP, and we find that AIRTP generally improve the quality scores.
@@
Appendix B Visualization of Creating Beam/Angle Plates

### v2

@@ Abstract
Radiotherapy (RT) planning is complex, subjective, and time-intensive. Advances in artificial intelligence (AI) promise to improve its precision, efficiency, and consistency, but progress is often limited by the scarcity of large, standardized datasets. To address this, we introduce the Automated Iterative RT Planning (AIRTP) system, a scalable solution for generating high-quality treatment plans. This scalable solution is designed to generate substantial volumes of consistently high-quality treatment plans, overcoming a key obstacle in the advancement of AI-driven RT planning. Our AIRTP pipeline adheres to clinical guidelines and automates essential steps, including organ-at-risk (OAR) contouring, helper structure creation, beam setup, optimization, and plan quality improvement, using AI integrated with RT planning software like Varian’s Eclipse. Furthermore, a novel approach for determining optimization parameters to reproduce 3D dose distributions, i.e. a method to convert dose predictions to deliverable treatment plans constrained by machine limitations is proposed. A comparative analysis of plan quality reveals that our automated pipeline produces treatment plans of quality comparable to those generated manually, which traditionally require several hours of labor per plan. Committed to public research, the first data release of our AIRTP pipeline includes nine cohorts covering head-and-neck and lung cancer sites to support an AAPM 2025 challenge. This data set features more than 10 times the number of plans compared to the largest existing well-curated public data set to our knowledge.
GitHub       Platform       Processed Data       DICOM Data       Pretrained Model
@@ 1 Introduction
Radiation therapy (RT) is a critical treatment modality used in approximately 50% of all cancer cases. Deep learning approaches have been incorporated into various stages of the RT planning process, including dose prediction [25, 6, 14], fluence prediction [34], leaf sequencing [13], and dose calculation [35]. Among recent advancements in RT, 3D dose prediction emerges as a particularly promising innovation. This approach directly estimates 3D dose distributions from inputs such as CT scans, planning target volume (PTV)/organ at risk (OAR) masks, and planning configurations, eliminating the need for labor-intensive optimization processes. Dose prediction serves multiple purposes, including supporting optimization objectives [6, 5], enabling quality assurance [16, 17], and functioning as a key component in fully AI-driven planning pipelines [13].
RT planning is a complex and subjective process involving multiple domain experts, making data curation from existing clinical plans highly labor-intensive due to human variability in contouring PTVs/OARs and planning practices. For instance, the number of OARs contoured and the quality of plans can vary significantly (see more in Sec. 2.1). Variations in radiotherapy planning introduce biases into training data, which are further exacerbated when data is collected from multiple cohorts. Additionally, most dose prediction studies face challenges such as a limited number of training patients and difficulties in quantitatively assessing quality metrics. Within the same institution, plan quality can also fluctuate over time due to the adoption of upgraded optimization technologies or updates to treatment protocols. These changes lead to inconsistencies in plan quality, further complicating the data used for AI training and evaluation. Those challenges leads to quality and quantity problems for AI training data, especially compare to well-known successfully AI models in other domain (extended discussion in Appendix A).
We see opportunities to automate the planning pipeline with high-quality. AI-based organ segmentation has achieved significant advancements in both research and commercial applications. Varian Medical Systems’ Eclipse, a widely used platform for optimizing radiotherapy treatment plans for cancer patients, supports automation through its Eclipse Scripting API (ESAPI). ESAPI enables developers to access and manipulate patient data for planning purposes, streamlining traditionally manual tasks performed via the Eclipse user interface. This functionality enhances workflow efficiency and ensures greater consistency in radiotherapy planning.
@@
  • We develop a method to replicate 3D dose distributions in both IMRT and VMAT scenarios using AIRTP, facilitating evaluation of dose prediction impacts on deliverable RT planning.
  • We have released the largest cleaned planning dataset to date for an AAPM challenge [^2] The GDP-HMM is one of the two challenges by the AAPM Working Group on Grand Challenges in 2025 (www.aapm.org/GrandChallenge). and beyond. This dataset: 1) contains over 10 times more plans than the existing curated public dataset [6], and 2) offers more practical settings to support diverse research explorations.
@@ 2.1 Challenges of Clinical Plans for AI Training
Additionally, some clinical plans were created more than five or ten years ago, when radiotherapy technology was less advanced. As a result, these plans may not meet the higher quality standards enabled by current technology. This highlights needs for re-planning RT treatments using the latest advancements while preserving clinical intents.
@@ 2.2 Automatic High Quality Planning at Scale
Step 2: Planning Script. There are three major modules in the planning script. Beam configuration includes setting up the fractions, beam geometries, and machine parameters. For each patient, we have both IMRT and VMAT setups following the same clinical goals. For the optimizer and dose calculation, we use TrueBeam with Photon Optimization version PO.18.0.1 and Dose Calculation AXB 18.0.1, respectively. To balance the time cost, the optimization convergence is set to “On” for VMAT.
@@
More detailed steps are outlined in Algorithm 1, along with some module implementations shown in Figures 10 and 11 (Appendix).
Input: Patient data with CT, PTV/OAR contours, helper structures, prescribed doses for PTVs; ScoreCard for plan evaluation; beam configurations; RapidPlan for a head-and-neck scenario. Output: A series of plans until the ScoreCard convergence. 1:Initialize plan based on RapidPlan or ScorePlan items 2:for $n$ = 1 to MAX iteration do 3:   Initialize objective buffer $D$ 4:   for $s$ in PTV and its Ring structures do 5:      Set the upper / lower point objectives based prescribed dose 6:      Use the same priorities as in the initial plan 7:   end for 8:   Calculate the ScoreCard for current plan 9:   for $s$ in ScoreCard OARs and Helpers do 10:      Calculate the DVH Cumulative Data $f_{s}(\cdot)$ 11:      for $v$ in volume points {1, 10, 20, 30, 40, 60, 80do 12:         Calculate the dose $d=f_{s}(v)$ 13:         Calculate the margin and objective dose and priority based on 10. 14:         Add the objective to objective buffer. 15:      end for 16:      for items in ScoreCard do 17:         Calculate the margin and objective dose and priority based on 11. 18:         Add the objective to objective buffer. 19:      end for 20:   end for 21:   optimize with the current objective buffer. 22:   update current plan. 23:end for
@@
(a) iterative high-quality planning
(b) translate 3D dose prediction to deliverable plan
3 Planning and Results
3.1 Planning Environment
3.2 Plan Quality Analysis
Fig. 4 demonstrates plan examples for lung cancer and head-and-neck cancer patients. Our AIRTP can generate arbitrarily large number of plans with varying parameters, given sufficient time and computing resources. As the first public release of our data, we include two plans per patient for the lung site and three plans per patient for the head-and-neck site. Additional examples can be found in Appendices F and G.
(a) Score improvement of IMRT plans.
(b) Score improvement of VMAT plans.
Figure 5: Quality improvement achieved through iterative refinement across head-and-neck cohorts.
3.3 Comparison with Clinical Plans
3.4 Plan Reproducibility Given 3D Dose
(a) HaN: reference (solid) vs. reproduce (dash).
(b) Lung: reference (solid) vs. reproduce (dash).
3.5 Beyond the Topics of Generalizable Dose Prediction
3.6 AI benchmarking for Generalizable Dose Prediction
Table 7: Data splits (train/valid/test).
# Plans # Patients
HaN 1428 / 224 / 298 499 / 77 / 100
Lung 1450 / 132 / 200 776 / 70 / 100
Total 3732 1622
Table 8: Results of different phases.
MAE ($\downarrow$) Plan Quality Score ($\uparrow$)
Sanity Check 1.981 103.09
Validation 2.540 110.76
Testing 3.454 118.78
Many backbones used in dose prediction have adopted architectures originally designed for medical image segmentation [14]. In the third-party medical image segmentation benchmark TouchStone [7], which includes a test set of 6,933 CT volumes, and the nnUNet revisit benchmark [22], which evaluates six public datasets, MedNeXt [30] achieves top performance, significantly outperforming popular Transformer-based backbones by a large margin (up to a 25% DSC improvement). Accordingly, we adopt MedNeXt as an educational baseline for generalizable dose prediction using publicly available data from our AIRTP, as illustrated in Fig. 13.
Inspired by [14], we introduce the angle plate and beam plate to encode spatial information related to beam angles. Additionally, categorical information (e.g., treatment mode, anatomical site) is incorporated as an extra input channel. The training dataset is divided into dev-train and dev-valid, where model weights are updated using dev-train, and the best checkpoint is selected based on dev-valid.
To facilitate participation, we provide a complete implementation—including code, data, tutorials, and a pre-trained model—on GitHub, enabling users with minimal radiotherapy (RT) background to easily engage with the challenge. An overview of the data distribution is presented in Table 8, while the baseline results are summarized in Table 8.333The plan quality scores for validation / testing phases are based on a randomly selected subset of 50 plans.
4 Discussion
As an initial step toward automating high-quality planning at scale, our work has several limitations compared to expert planning in clinical practice. First, clinical planning involves detailed patient-specific considerations, such as variations in the contouring of PTVs/OARs, auxiliary structures, and various dose constraints. In our pipeline, we focus primarily on clinical PTVs and prescribed doses as the key patient-specific factors, simplifying the process due to the lack of access to more detailed patient data. Second, different institutions often follow different protocols, and even within the same institution, planners may have individual preferences. Our pipeline does not currently account for such subjective variations. While these preferences are partially reflected in the scorecards, we use only one standardized scorecard per treatment site. A future direction is to collaborate with dosimetrists to develop multiple scorecards that better capture diverse clinical preferences and practices. Third, unlike more well-established research fields, our comparative analyses of iterative refinement and re-planning strategies are limited by the lack of directly relevant literature. Although we made efforts to compare the DVHs of AIRTP plans with clinically approved plans, this comparison may not be entirely fair, as the AIRTP and clinical plans follow different guidelines. Fourth, this paper focuses on 1) novel methodology to automate high-quality planning at scale, 2) analyses of planning results, 3) open sourcing high-quality data/tutorial/code/pre-trained model for a AAPM challenge and other RT research topics. the summary and clinical impact analysis of the GDP-HMM challenge are beyond the scope of this paper and will be provided following the conclusion of the challenge.
Appendix
[…]
Appendix A Data Scale of Popular Successfully AI Models and RT Planning Studies
Figure 9: An illustration of what a huge gap between AI models in RT planning and some well-known successful AI models in other domain. The y-axis is in base-10 logarithmic scale.
We add more descriptions about the models in Fig. 9:
  • Llama 3 [15]: An advanced large language model (LLM) designed for natural language understanding and generation, offering improved efficiency and performance over its predecessors.
  • Stable Diffusion [29]: A state-of-the-art latent diffusion model for generating high-quality images from text prompts, widely used for AI-driven image synthesis.
  • Segment Anything Models (SAM) [24]: A foundational vision model capable of segmenting objects in images with minimal user input, enabling versatile and efficient image annotation.
  • Medical SAM (Med-SAM) [27]: A specialized adaptation of SAM for medical imaging, fine-tuned to segment anatomical structures with high accuracy across diverse modalities.
  • RETFound [36]: A foundation model for retinal imaging that leverages self-supervised learning to enhance disease detection and diagnosis in ophthalmology.
  • Medical AI for Synthetic Imaging (MAISI) [19]: A generative AI framework designed to synthesize medical images for training and validation, improving AI model generalization and robustness.
  • Dose Prediction Models [14, 6, 12, 23]: AI-based models for predicting radiation therapy dose distributions, as the task in GDP-HMM challenge.
  • Fluence Prediction Models [33, 26]: Computational models that predict fluence maps in radiation therapy, helping to optimize beam intensity for effective dose delivery.
  • Leaf Sequencing Models [13, 21]: Algorithms that translate optimized fluence maps into deliverable multi-leaf collimator (MLC) sequences, ensuring efficient and accurate radiation therapy delivery.
Many of the well-known successful AI models have been trained with large-scale and well-curated data. In contrast, there are significant challenges for AI application in RT: 1) AI training samples in the RT planning domain are rather limited, usally under 1,000 sample; 2) missing and noisy information usually exists during clinical RT data collection; and 3) most well-curated RT training samples are not shared publicly. With the enthusiasm to advance RT planning with AI, we conducted this study to address those challenges.
Appendix B Code Blocks for Algorithm 1
@@
Appendix C More about GDP-HMM Challenge
C.1 Data Distribution
Table 9: Data Distribution for GDP-HMM Challenge.
Training Split Validation Split Test Split
#Plans
HaN 1426 (49.6%) 224 (62.9%) 298 (59.8%)
Lung 1450 (50.4%) 132 (37.1%) 200 (40.2%)
#Patients
HaN 498 (39.1%) 77 (52.4%) 100 (50.0%)
Lung 776 (60.9%) 70 (47.6%) 100 (50.0%)
#Plan by #PTV
3PTVs per Plan (HaN) 912 (31.7%) 125 (35.1%) 148 (29.7%)
2PTVs per Plan (HaN) 413 (14.4%) 84 (23.6%) 120 (24.1%)
1PTV per Plan (HaN) 101 (3.5%) 15 (4.2%) 30 (6.0%)
1PTV per Plan (Lung) 1450 (50.4%) 132 (37.1%) 200 (40.2%)
#Plans by Type
IMRT 1645 (57.2%) 212 (59.6%) 300 (60.2%)
VMAT 1231 (42.8%) 144 (40.4%) 198 (39.8%)
Figure 12: Kernel density estimation of the high prescription dose in different data splits.
The data distribution of GDP-HMM challenge is shown in Table 9 and Fig. 12. The testing patients are from different institutes than the training and validation splits.
C.2 Baseline Method
As shown in Fig. 13, the educational baseline applies the MedNeXt as the backbone; due to its superior performance in multiple large-scale segmentation benchmarks. Fundamentally, the dose prediction is an conditional image generation problem. The CT images, PTV & OAR masks, planning mode, treatment site, and beam geometries etc., are conditions for 3D dose prediction. We follow the strategy of from [14], concatenating all the conditions as multi-channel 3D input. The loss function used in the baseline is pretty simple: the mean absolute error (MAE) of the predicted dose and the reference dose.
For more details, please refer to links of code & tutorial and pre-trained model.
Figure 13: A baseline for the generalizable dose prediction for GDP-HMM challenge. The MedNeXt is used as backbone (U-shape Network).
C.3 Result Samples from the Baseline
Figure 14: Head-and-neck IMRT example.
Figure 15: Head-and-neck VMAT example.
We present four samples from our validation cohort, illustrating distinct scenarios that a universal AI model should be capable of handling. Figures 14 and 15 depict treatment plans for the same head-and-neck cancer patient using different modalities: IMRT and VMAT. Similarly, Figures 16 and 17 show treatment plans for the same lung cancer patient, also with IMRT and VMAT. Each case includes a comparison of the predicted and reference 3D dose distributions, along with corresponding dose-volume histograms (DVHs).
Figure 16: Lung IMRT example.
Figure 17: Lung VMAT example.
Appendix D Scope of This Paper and the Planned Challenge Summary Paper
A summary paper for the challenge is planned to be published several months after its conclusion. It is important to highlight the distinctions between this paper and the upcoming challenge summary paper.
This paper presents our key contributions, including: 1) A novel AIRTP pipeline for automatically generating high-quality radiotherapy plans. 2) A comprehensive dataset (DICOM and NPZ format) and accompanying toolkit to facilitate research. 3) Baselines and tutorials for the challenge.
The advantages of AIRTP, outlined in Section 1, enable large-scale AI training in radiotherapy, addressing the limitations of previous studies, which often relied on fewer than 1,000 plans. Our pipeline streamlines data generation, allowing for significantly larger training scales.
Additionally, our dataset supports a broad range of research topics beyond dose prediction, the primary focus of the challenge. We also provide a toolkit for reading and visualizing raw DICOM data, enhancing usability for researchers.
Furthermore, we establish a benchmark for the GDP-HMM challenge using the state-of-the-art MedNeXt model. We provide detailed tutorials, open-source code, pre-trained model weights, and thorough result analyses, offering participants and future researchers a strong foundation for further exploration.
The challenge itself is focused solely on dose prediction, which is just one of many critical topics in radiotherapy. The planned challenge summary paper will primarily cover: 1) A comparative analysis of participants’ methods, evaluating AI techniques and performance results. 2) A deeper discussion on how AI can enhance radiotherapy planning through 3D dose prediction, with a focus on clinical implications.
This planned summary will provide insights into the practical impact of AI in radiotherapy, complementing the technical contributions outlined in this paper.
Appendix E Detailed Scorecard Examples
@@
Table 10 and 11 shows the details scores and final score from the scorecard [32] to evaluate plan qualities for IMRT and VMAT, respectively. The structures (from the automatic pipeline) too small or with too much overlap with PTVs have been removed. For each table, we have included the RT structure names (i.e., StructId), Metric Type, the input for the Metric Type, output value, and the score of the output. We compare the output and its value of the plans from the initial (without the iterative refining) and full AIRTP, and we find that AIRTP generally improve the quality scores.
@@
Appendix F Visualization of Creating Beam/Angle Plates